# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v2

@@ Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances exhibiting complex geometries or significant deviations from canonical shapes. To address this challenge, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our approach first predicts semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Generator, then refines them with: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequence. Additionally, we design a surface loss and a separation loss to enforce uniform coverage and spatial diversity in keypoint distribution. The generated keypoints are finally mapped to a canonical space for regressing the object’s 6D pose and size. Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance and significantly outperforms existing methods.
@@ I Introduction
Object pose estimation is a fundamental yet challenging problem in computer vision and robotics, with broad applications in robotic grasping [1], augmented reality [2], 3D scene understanding [3], and autonomous driving [4]. Existing approaches can be broadly classified into instance-level [5, 6, 7, 8, 9] and category-level [10, 11, 12, 13, 14] methods. Instance-level methods rely on specific CAD models to estimate the poses of known objects. However, their heavy dependence on accurate CAD models makes them difficult to generalize to real-world scenarios, where objects are often novel and lack corresponding models. This severely limits their applicability to unseen instances within the same category or to objects with variations in shape and appearance. As a more flexible alternative, category-level object pose estimation aims to predict the pose and size of previously unseen instances from predefined object categories. Despite its potential, this task remains considerably more challenging due to the significant intra-class variations in object shape, texture, and structure.
@@
Among recent advances, correspondence-based methods [10, 11, 15, 16, 13] have demonstrated promising performance in category-level object pose estimation. These approaches typically rely on establishing dense 3D-3D correspondences between observed point clouds and category-level canonical spaces, such as the Normalized Object Coordinate Space (NOCS)[10]. In NOCS, all instances from the same category are mapped into a unified 3D coordinate frame, and the object pose and size are recovered by estimating a similarity transformation between the observed object and its canonical reconstruction using the Umeyama algorithm [17]. As a result, the quality of the reconstructed NOCS representation directly influences the final pose estimation accuracy.
To improve the quality of NOCS reconstructions, a common approach is to incorporate category-level shape priors. For example, SPD [11] learns object-specific deformations from predefined category templates and establishes dense correspondences between the observed point cloud and its reconstructed canonical shape. This paradigm has inspired a series of follow-up works [15, 16, 18, 19], which typically employ average shape point clouds as priors and design dedicated deformation networks to better adapt to intra-class variations. However, such priors remain static during both training and inference, limiting the model’s adaptability to diverse and unseen geometries. To address this, Query6DoF [20] introduces class-specific learnable query embeddings that are jointly optimized with the pose estimator, allowing the priors to adapt dynamically during training. More recently, shape-prior-free methods [12, 21, 13, 22, 23, 14] have drawn increasing attention due to their enhanced generalization to novel object instances. Among them, AG-Pose [14] introduces a keypoint-based method that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving notable improvements against structural discrepancies. Nevertheless, its keypoint features are derived solely from local attention within fixed neighborhoods and simple global average pooling, limiting their capacity to capture fine-grained geometric details and hindering the model’s ability to perceive global structural consistency. Moreover, the keypoint distribution is regulated by heuristic losses, lacking explicit constraints to ensure uniform surface coverage and sufficient spatial separation. As a result, the predicted keypoints (as illustrated in Fig. 1) often cluster in limited regions, fail to accurately capture the object’s surface coverage, and exhibit substantial NOCS errors, particularly on objects with complex geometries.
@@
To address the above challenges, we propose INKL-Pose, a novel category-level object pose estimation method that enables instance-adaptive keypoint learning with local-to-global geometric aggregation. Specifically, we first extract and fuse RGB and point cloud features using a Pointwise Feature Encoder, and then generate a sparse set of semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Generator. These keypoints are subsequently refined by two complementary modules: (1) a Local Keypoint Feature Aggregator that captures fine-grained geometric details from spatial neighborhoods, and (2) a Global Keypoint Feature Aggregator that utilizes a bidirectional Mamba architecture to model long-range dependencies while maintaining structural consistency across the object. To handle the unordered and non-causal nature of point clouds, we introduce a Feature Sequence Flipping strategy that facilitates bidirectional perception in the feature space while preserving the original spatial structure. Additionally, we design a surface loss and a separation loss to explicitly encourage uniform coverage and spatial diversity among keypoints. The refined keypoints are then mapped to a canonical space to regress the object’s 6D pose and size. As illustrated in Fig. 1, even on complex camera instances, the keypoints predicted by our INKL-Pose still achieve accurate and consistent surface coverage, and exhibit significantly lower NOCS errors compared to those predicted by AG-Pose[14].
@@ II-A Instance-Level Object Pose Estimation
Instance-level object pose estimation aims to infer the precise pose of an object based on its corresponding 3D CAD model. Existing approaches in this domain can be broadly divided into three categories: direct regression methods [6, 24], direct voting methods [25, 26] and dense 2D-3D correspondence methods [27, 9]. Direct regression methods utilize end-to-end networks to directly predict the object’s rotation and translation parameters from extracted visual features. Although this approach simplifies the pose estimation pipeline, it suffers from reduced accuracy in cluttered or occluded scenes due to the limitations of direct mapping. Direct voting methods, on the other hand, estimate pose candidates and their corresponding confidence scores at the pixel or point level, ultimately selecting the pose with the highest confidence. While these methods are capable of capturing detailed local information and generally yield strong performance, the dense voting process can be computationally intensive and time-consuming. Alternatively, the third category of instance-level methods adopts a correspondence-based paradigm, which first establishes dense 2D–3D correspondences between RGB images and the complete 3D CAD model, and subsequently estimates the object pose using algorithms such as Perspective-n-Point or least-squares optimization. Unlike direct regression or voting-based approaches, this method explicitly decouples feature matching from pose estimation, thereby enabling more geometry-aware inference when accurate correspondences are available. However, the performance of such methods heavily depends on the availability of precise CAD models and robust feature descriptors, both of which are often difficult to obtain in real-world scenarios.
@@ II-B Category-Level Object Pose Estimation
Unlike instance-level approaches that rely on known CAD models, category-level 6D object pose estimation aims to estimate the pose and size of previously unseen instances from predefined object categories. This setting is inherently more general and challenging due to substantial intra-class variations among object instances. To address these challenges, NOCS [10] introduces the Normalized Object Coordinate Space, which maps different instances to a unified canonical frame, thereby mitigating variations in scale and geometry. The Umeyama algorithm [17] is then applied to recover the object pose via similarity transformation. To better handle intra-class variations, SPD [11] proposes a deformation network that reconstructs object geometry by learning instance-specific deformations from category-level shape priors, followed by dense correspondence matching between the reconstructed canonical model and the observed point cloud. This paradigm has inspired a series of follow-up works [18, 19, 15, 16], which further explore the extraction of more robust features and the development of more efficient deformation strategies. However, the static nature of shape priors during both training and inference limits adaptability and hinders generalization to objects with substantial structural variations.
To overcome this limitation, Query6DoF [20] introduces class-specific learnable queries that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [12, 21, 13, 22, 23, 14] have attracted increasing attention for their improved generalization to diverse and unseen object geometries. IST-Net [13] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [12] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. TG-Pose [21] introduces topological descriptors and a symmetry-aware reconstruction module to refine object geometry and enhance pose stability. SecondPose [23] integrates geometric features with semantic embeddings from DINOv2 [28], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [29] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. CleanPose [30] integrates causal learning with knowledge distillation to address the spurious correlations introduced by confounding factors in pose estimation. More recently, AG-Pose [14] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
@@ II-C State Space Models
Motivated by these advances, recent works [34, 35, 36, 37, 38] have adapted the Mamba architecture for 3D point cloud analysis. For instance, PointMamba [34] transforms unordered point cloud data into ordered 1D sequences using Hilbert and Trans-Hilbert curves, preserving spatial continuity and semantic relevance. PCM [35] further utilizes bidirectional Mamba with a point sequence flipping strategy to allow each point to access features from any other point in the cloud. Inspired by the success of Mamba in long-range sequence modeling, we explore its potential for category-level object pose estimation. In particular, we leverage bidirectional Mamba to model globally consistent geometric structures across varied object instances, and propose a novel bidirectional input sequence generation strategy tailored to the unordered and non-causal nature of point cloud data.
@@ III-A Overview
In this section, we present the detailed architecture of the proposed INKL-Pose. As illustrated in Fig. 1(a), our objective is to jointly estimate an object’s 3D rotation $R\in\mathbb{R}^{3\times 3}$, 3D translation $t\in\mathbb{R}^{3}$ and size $s\in\mathbb{R}^{3}$ from an input RGB image $\mathcal{I}_{o}$ and its corresponding point cloud $\mathcal{P}_{o}$. Specifically, INKL-Pose consists of five key stages. First, the Pointwise Feature Encoder (Section III-B) extracts instance features from both the RGB image and the point cloud, which are subsequently fused to facilitate cross-modal integration. Second, the fused features are processed by the Instance-Adaptive Keypoint Generator (Section III-C) to predict a sparse set of keypoints and their associated features. Third, the Local Keypoint Feature Aggregator (Section III-D) enriches each keypoint with geometrically distinct local features. Fourth, the Global Keypoint Feature Aggregator (Section III-E) captures globally consistent geometric context across object instances. Finally, the Pose and Size Estimator (Section III-F) takes the enhanced keypoints as input and directly regresses the object’s 6D pose and size.
Figure 2: (a) Overall framework of INKL-Pose [^1] This figure was inspired by the drawing style of CleanPose [30]. We gratefully acknowledge the authors for their work.. (b) The Instance-Adaptive Keypoint Generator dynamically generates keypoints along with corresponding features by assigning weights based on the semantic and geometric information extracted by the Pointwise Feature Encoder. (c) The Local Keypoint Feature Aggregator captures fine-grained local structures by aggregating point cloud neighborhood information. (d) The Global Keypoint Feature Aggregator models long-range dependencies using bidirectional Mamba, extracting globally consistent geometric representations across instances.
@@ III-B Pointwise Feature Encoder
III-C Instance-Adaptive Keypoint Generator
Given the challenges of noise and incomplete observations in cluttered environments, we propose an Instance-Adaptive Keypoint Generator (IAKG) that dynamically generates keypoints from the most informative regions of each object, ensuring robustness and adaptability across diverse instances. Rather than relying on fixed strategies like Farthest Point Sampling (FPS), which select keypoints solely based on spatial distances and fail to adapt to the geometric diversity across instances, our method learns to predict instance-specific keypoints in a data-driven and context-aware manner.
Specifically, we formulate keypoint generation as a learnable query-based process that adapts to both semantic and geometric variations, following the paradigm introduced in AG-Pose [14]. In contrast to AG-Pose, our approach introduces two novel loss functions that explicitly encourage the predicted keypoints to be uniformly distributed and spatially diverse across the object surface. As shown in Fig. 1(b), we first initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$, which is initially set to zero, providing a neutral starting point that avoids introducing bias toward specific patterns. This initialization facilitates stable learning dynamics and allows the query vectors to adapt purely based on the input data through the attention mechanism. Then, through multi-head cross-attention and multi-head self-attention mechanisms, the keypoint query vectors interact comprehensively with both the input features $\mathcal{F}_{o}$ and themselves, progressively updating to form instance-specific query embeddings. Finally, we compute the cosine similarity between $\mathcal{Q}_{kpt}$ and $\mathcal{F}_{o}$, and apply a softmax function to normalize the results, yielding the keypoint probability distribution matrix $\mathcal{W}$:
$\mathcal{W}=Softmax\left(\frac{\mathcal{Q}_{kpt}\cdot\mathcal{F}_{o}}{\|\mathcal{Q}_{kpt}\|_{2}\cdot\|\mathcal{F}_{o}\|_{2}+\varphi}\right),$ (1)
@@ III-C Instance-Adaptive Keypoint Generator
where $\text{NN}_{j}(\mathcal{P}_{kpt}^{i})$ denotes the $j$th nearest neighbor of the $i$th keypoint, and $M$ is the number of neighbors considered. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint to mitigate keypoint clustering while maintaining flexibility in spatial layout. The denominator is lower-bounded by a small constant $\epsilon$ = 0.01 to ensure numerical stability.

### v3

@@ Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances exhibiting complex geometries or significant deviations from canonical shapes. To address this issue, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our method first predicts semantically consistent and geometrically informative keypoints using an Instance-Adaptive Keypoint Detector, then refines them: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a simple yet effective Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequence. Additionally, we design a surface loss and a separation loss to encourage uniform coverage and spatial diversity in keypoint distribution. The resulting keypoints are mapped to a canonical space for 6D pose and size regression. Extensive experiments on CAMERA25, REAL275, and HouseCat6D show that INKL-Pose achieves state-of-the-art performance with 16.7M parameters and runs at 36 FPS on an NVIDIA RTX 4090D GPU.
@@ I Introduction
Object pose estimation is a fundamental yet challenging problem in computer vision and robotics, with broad applications in robotic grasping [1], autonomous driving [2], and 3D scene understanding [3]. Existing approaches can be broadly classified into instance-level [4, 5, 6, 7, 8, 9] and category-level [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] methods. Instance-level methods rely on specific CAD models to estimate the poses of known objects. However, their heavy dependence on accurate CAD models makes them difficult to generalize to real-world scenarios, where objects are often novel and lack corresponding models. This severely limits their applicability to unseen instances within the same category or to objects with variations in shape and appearance. As a more flexible alternative, category-level object pose estimation aims to predict the pose and size of previously unseen instances from predefined object categories. Despite its potential, this task remains considerably more challenging due to the significant intra-class variations in object shape, texture, and structure.
@@
Recent advances in correspondence-based methods [10, 11, 12, 13, 14] have shown promising performance in category-level object pose estimation. However, these approaches typically rely on static priors that restrict their adaptability to novel object instances. To address this limitation, shape-prior-free methods [15, 16, 17, 18, 19, 20] have drawn increasing attention due to their enhanced generalization to novel object instances. Among these methods, AG-Pose [20] proposes a keypoint-based framework that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving significant improvements in handling structural discrepancies. Despite its success, AG-Pose derives keypoint features solely from local attention within fixed neighborhoods, combined with a simple global average pooling strategy. This design limits its capacity to capture fine-grained geometric cues and to model global structural consistency. Although AG-Pose incorporates losses to encourage uniformity and separation in keypoint distribution, their effectiveness diminishes when dealing with objects exhibiting complex or irregular geometries. As illustrated in Fig. 1, the resulting keypoints often cluster in limited regions, leading to incomplete surface representation and substantial NOCS errors.
@@
To address the above challenges, we propose INKL-Pose, a novel category-level object pose estimation method that enables instance-adaptive keypoint learning with local-to-global geometric aggregation. Specifically, we first extract and fuse RGB and point cloud features using a Pointwise Feature Encoder, and then generate a sparse set of semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Detector. These keypoints are subsequently refined by two complementary modules: (1) a Local Keypoint Feature Aggregator that captures fine-grained geometric details from spatial neighborhoods, and (2) a Global Keypoint Feature Aggregator that utilizes a bidirectional Mamba architecture to model long-range dependencies while maintaining structural consistency across the object. To handle the unordered and non-causal nature of point clouds, we introduce a simple yet effective Feature Sequence Flipping strategy that facilitates bidirectional perception in the feature space while preserving the original spatial structure. Additionally, we design a surface loss and a separation loss to explicitly encourage uniform coverage and spatial diversity among keypoints. The refined keypoints are then mapped to a canonical space to regress the object’s 6D pose and size. As illustrated in Fig. 1, even on complex camera instances, the keypoints predicted by our INKL-Pose still achieve accurate and consistent surface coverage, and exhibit significantly lower NOCS errors compared to those predicted by AG-Pose [20].
@@ II-A Instance-Level Object Pose Estimation
Instance-level object pose estimation aims to infer the precise pose of an object based on its corresponding 3D CAD model. Existing approaches in this domain can be broadly divided into three categories: direct regression methods [4, 5], direct voting methods [6, 7] and dense 2D-3D correspondence methods [8, 9]. Direct regression methods utilize end-to-end networks to directly predict the object’s rotation and translation parameters from extracted visual features. Although this approach simplifies the pose estimation pipeline, it suffers from reduced accuracy in cluttered or occluded scenes due to the limitations of direct mapping. Direct voting methods, on the other hand, estimate pose candidates and their corresponding confidence scores at the pixel or point level, ultimately selecting the pose with the highest confidence. While these methods are capable of capturing detailed local information and generally yield strong performance, the dense voting process can be computationally intensive and time-consuming. Alternatively, the third category of instance-level methods follows a correspondence-based paradigm, which first establishes dense 2D–3D correspondences between RGB images and their associated 3D CAD models, and then estimates the object pose using algorithms such as Perspective-n-Point (PnP) or least-squares optimization. Unlike direct regression or voting-based approaches, this method explicitly decouples feature matching from pose estimation, enabling more reliable, geometry-informed inference when accurate correspondences are available. However, their performance heavily relies on the availability of precise CAD models and robust feature descriptors, both of which are often difficult to obtain in real-world scenarios.
@@ II-B Category-Level Object Pose Estimation
Unlike instance-level approaches that rely on known CAD models, category-level 6D object pose estimation aims to estimate the pose and size of previously unseen instances from predefined object categories. This setting is inherently more general and challenging due to substantial intra-class variations among object instances. To address these challenges, NOCS [10] introduces the Normalized Object Coordinate Space, which maps different instances to a unified canonical frame, thereby mitigating variations in scale and geometry. The Umeyama algorithm [22] is then applied to recover the object pose via similarity transformation. To better handle intra-class variations, SPD [11] proposes a deformation network that reconstructs object geometry by learning instance-specific deformations from category-level shape priors, followed by dense correspondence matching between the reconstructed canonical model and the observed point cloud. This paradigm has inspired a series of follow-up works [12, 13, 14], which further explore the extraction of more robust features and the development of more efficient deformation strategies. However, the static nature of shape priors during both training and inference limits adaptability and hinders generalization to objects with substantial structural variations.
To overcome this limitation, Query6DoF [15] introduces class-specific learnable queries that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [16, 17, 18, 19, 20] have attracted increasing attention for their improved generalization to diverse and unseen object geometries. IST-Net [17] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [16] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. SecondPose [19] integrates geometric features with semantic embeddings from DINOv2 [23], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [21] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. CleanPose [24] integrates causal learning with knowledge distillation to address the spurious correlations introduced by confounding factors in pose estimation. More recently, AG-Pose [20] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
@@ II-C State Space Models
Motivated by these advances, recent studies have adapted the Mamba architecture for 3D point cloud analysis. For instance, PointMamba [28] transforms unordered point cloud data into ordered 1D sequences using Hilbert and Trans-Hilbert curves, preserving spatial continuity and semantic relevance. PCM [29] further utilizes bidirectional Mamba with a point sequence flipping strategy to allow each point to access features from any other point in the cloud. Inspired by the success of Mamba in long-range sequence modeling, we explore its potential for category-level object pose estimation. In particular, we leverage bidirectional Mamba to model globally consistent geometric structures across varied object instances, and propose a novel bidirectional input sequence generation strategy tailored to the unordered and non-causal nature of point cloud data.
@@ III-A Overview
In this section, we present the detailed architecture of the proposed INKL-Pose. As illustrated in Fig. 2, our objective is to jointly estimate an object’s 3D rotation $R\in\mathbb{R}^{3\times 3}$, 3D translation $t\in\mathbb{R}^{3}$ and size $s\in\mathbb{R}^{3}$ from an input RGB image $\mathcal{I}_{o}$ and its corresponding point cloud $\mathcal{P}_{o}$. Specifically, INKL-Pose consists of five key stages. First, the Pointwise Feature Encoder (Section III-B) extracts instance features from both the RGB image and the point cloud, which are subsequently fused to facilitate cross-modal integration. Second, the fused features are processed by the Instance-Adaptive Keypoint Detector (Section III-C) to predict a sparse set of keypoints and their associated features. Third, the Local Keypoint Feature Aggregator (Section III-D) enriches each keypoint with geometrically distinct local features. Fourth, the Global Keypoint Feature Aggregator (Section III-E) captures globally consistent geometric context across object instances. Finally, the Pose and Size Estimator (Section III-F) takes the enhanced keypoints as input and directly regresses the object’s 6D pose and size.
Figure 2: Overall framework of INKL-Pose. The Instance-Adaptive Keypoint Detector dynamically generates keypoints along with corresponding features using semantic and geometric information extracted by the Pointwise Feature Encoder. The Local Keypoint Feature Aggregator captures fine-grained local structures by aggregating neighborhood information from the point cloud. The Global Keypoint Feature Aggregator models long-range dependencies using bidirectional Mamba, extracting globally consistent geometric representations across instances. The Pose and Size Estimator predicts object rotation, translation, and size from the keypoints and their features.
@@ III-B Pointwise Feature Encoder
III-C Instance-Adaptive Keypoint Detector
Given the challenges of noise and incomplete observations in cluttered environments, we adopt the Instance-Adaptive Keypoint Detector (IAKD) from AG-Pose [20] as our keypoint generation backbone. This detector formulates keypoint generation as a learnable, query-based process that dynamically generates keypoints from the most informative regions of each object. Specifically, we initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$ set to zero, providing a neutral starting point that avoids bias toward specific patterns. Through multi-head cross-attention and self-attention mechanisms, these query embeddings iteratively interact with the input features $\mathcal{F}_{o}$ and themselves, progressively refining to capture instance-specific keypoint features. The keypoint probability distribution matrix $\mathcal{W}$ is computed using cosine similarity followed by a softmax function:
$\mathcal{W}=\text{Softmax}\left(\frac{\mathcal{Q}_{kpt}\cdot\mathcal{F}_{o}}{\|\mathcal{Q}_{kpt}\|_{2}\cdot\|\mathcal{F}_{o}\|_{2}+\varphi}\right),$ (1)
@@ III-C Instance-Adaptive Keypoint Detector
where $\text{NN}_{j}(\mathcal{P}_{kpt}^{i})$ denotes the $j$-th nearest neighbor of the $i$-th keypoint, and $M$ is the number of neighbors considered. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint to mitigate keypoint clustering while maintaining flexibility in spatial layout. The denominator is lower-bounded by a small constant $\epsilon$ = 0.01 to ensure numerical stability.
@@
Figure 3: Illustration of the Local Keypoint Feature Aggregator (LKFA) and Global Keypoint Feature Aggregator (GKFA). (a) LKFA leverages KNN-based neighborhood construction and positional encoding, followed by feature expansion and aggregation, to capture fine-grained geometric details from local neighborhoods. (b) GKFA applies feature sequence flipping to obtain a backward sequence and leverages bidirectional Mamba to model long-range dependencies and global shape structures by integrating both original and flipped feature sequences.